2. (15 points) The table below gives the amount of output (Q) that can be produced with various combinations of capital (K) and labor (L

K L Q TFC TVC TC ATC MC

10 0 0

10 1 14

10 2 35

10 3 62

10 4 91

10 5 121

The price of labor is w=$100/unit and the price of capital is v=$50/unit. Complete the above cost schedule by determining total fixed cost (TFC), total variable cost (TVC), total cost (TC), average total cost (ATC) and marginal cost (MC).
Answer #1
K L Q TFC TVC TC ATC MC
10 0 0 500 0 500
10 1 14 500 100 600 42.86 7.14
10 2 35 500 200 700 20.00 4.76
10 3 62 500 300 800 12.90 3.70
10 4 91 500 400 900 9.89 3.45
10 5 121 500 500 1000 8.26 3.33

TFC=K*v(price of capital)

TVC=L*w(wage)

TC=TVC+TFC

ATC=TC/Q

MC=chnage in TC/chnage in Q